OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1))

jimthompson5910 (jim_thompson5910):

First find g(-1) \[\large g(x) = x^2 - 7x - 9\] \[\large g(-1) = (-1)^2 - 7(-1) - 9\] \[\large g(-1) = 1 - 7(-1) - 9\] \[\large g(-1) = 1 + 7 - 9\] \[\large g(-1) = 8 - 9\] \[\large g(-1) = -1\]

jimthompson5910 (jim_thompson5910):

then you use this value of g(-1) to find f(g(-1)) f(x) = x - 2 f(g(-1)) = g(-1) - 2 ... replace each x with g(-1) f(g(-1)) = -1 - 2 ... replace g(-1) on the right side with -1 (since g(-1) = -1) f(g(-1)) = -3

jimthompson5910 (jim_thompson5910):

coincidentally, g(-1) = -1, but the input and output are usually not the same

OpenStudy (anonymous):

so f(g(-1)) = -3? im starting to understand this lesson

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (anonymous):

i understood everything till f(x) came in the equation lol but thank you! im starting to understand it more

jimthompson5910 (jim_thompson5910):

well you start with f(x) = x-2 then you replace x with g(-1) since you want f(g(-1))

jimthompson5910 (jim_thompson5910):

so that explains how we got f(g(-1)) = g(-1) - 2 then on the right side you replace g(-1) with -1 because above we've shown that g(-1) = -1

OpenStudy (anonymous):

OHHHHHHHHHHHH I see

jimthompson5910 (jim_thompson5910):

glad it's clicking now

OpenStudy (anonymous):

is easy. its just that if you think it too much it gets confusing

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yeah just take it one step at a time and you'll be fine

OpenStudy (anonymous):

i know! thank you! but what about ? Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is g(2)?

OpenStudy (anonymous):

g(2) = x2 -6x - 7 g(2) = 2^2 - 6x - 7 ???

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need to replace the x in -6x as well

jimthompson5910 (jim_thompson5910):

g(x) = x^2 - 6x - 7 g(2) = (2)^2 - 6(2) - 7 g(2) = ???

OpenStudy (anonymous):

yes yes i know that i forgot to type it in tho lol. g(2) = -15?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now that you know that g(2) = -15, you basically plug -15 into f(x) and evaluate

OpenStudy (anonymous):

im learning better than in class lol. ok let me work that out

jimthompson5910 (jim_thompson5910):

this is because f(g(2)) = f(-15) effectively since g(2) = -15

jimthompson5910 (jim_thompson5910):

that's great you are

OpenStudy (anonymous):

so.... f(x) = x + 8 f(-15) = -15 + 8 f(-15) = -7?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so that means f(g(2)) = -7

OpenStudy (anonymous):

wait so the final answer is f(g(2)) = -7? THANK YOU

jimthompson5910 (jim_thompson5910):

yep, it is

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw